# Dust destruction by the supernova remnant forward shock in a turbulent interstellar medium

### v1

@@ Abstract
Context. While supernova remnants (SNRs) are observed to produce up to 1 M of dust, the amount of dust destroyed by the forward shock is poorly constrained, raising the question whether they are net dust producers or destroyers.
Aims. We aim to estimate the dust destruction efficiency of SNR forward shocks in a realistically turbulent interstellar medium (ISM) during their most destructive phase, and assess dust shielding by high density filaments during this period.
Methods. We run 3D high-resolution turbulence simulations for different turbulent Mach numbers (0–3) and average ISM densities (1–100 cm-3) to resemble observations of the turbulent ISM. We then set off a supernova explosion to trace its 3D magnetohydrodynamical evolution for the first 10 kyr. Finally, we run post-processing simulations to investigate the dust transport and destruction by the SNR forward shock, taking into account gas and plasma drag, kinetic and thermal sputtering, and grain-grain collisions, and considering either silicates or carbonaceous dust.
Results. The dust destruction rate of the forward shock strongly depends on the average ISM density and turbulence strength, varying between 27–92% (0.85–11.0 M) in the studied 10 kyr. Overall, dust is less efficiently destroyed in a low density medium ($1\,\mathrm{cm^{-3}}$, 27–57%) than in intermediate ($10\,\mathrm{cm^{-3}}$, 46–92%) and high density ($100\,\mathrm{cm^{-3}}$, 73–87%) cases. The forward shock is found to destroy 8–34% less dust in high Mach turbulence compared to a homogeneous medium. Furthermore, carbonaceous grains are more robust (up to 21% more) than silicates.
Conclusions. Filaments can partly shield dust from destruction in the first 10 kyr, however, always more than 0.85 M of dust is destroyed, making most SNRs dust sinks under the conditions explored in this work. The destruction efficiency of the SNRs with less than 1 M of destroyed dust has not yet plateaued so that they are most likely also net dust destroyers by the end of their lifetime.
@@ 1 Introduction
Understanding dust formation and destruction is key for planetary, stellar, and (extra-)galactic astronomy alike (c.f. Draine2003; Micelotta2018). Asymptotic giant branch (AGB) stars and supernova remnants (SNRs) are known to provide favorable conditions for dust production. Formed dust grains are also hypothesized to be able to grow efficiently in the interstellar medium (ISM; e.g. Draine2009; Hirashita2012; Mattsson2012; Zhukovska2016; Zhukovska2018), significantly increasing the ISM dust budget over time. Yet, the relative importance of these dust formation and growth mechanisms across redshifts and the detailed underlying physical processes remain uncertain (c.f. Schneider2024).
@@
AGB stars play an important role in the dust budget of local galaxies (e.g. Matsuura2009; Olofsson2010; Schneider2014; Dell'Agli2015). However, at high redshifts, where several galaxies have been observed to contain high dust masses (e.g. Pettini1994; Bertoldi2003; Watson2015; Bowler2022; Sommovigo2022; Akins2023; Witstok2023), the dust needs to be produced at a much faster rate than typical AGB dust production channels can provide. AGB stars are end evolutionary stages of low and intermediate mass stars with masses of less than 8 M and, thus, take several tens of million years to start contributing to the dust budget (e.g. Vassiliadis1993; Valiante2009). If the initial mass function is bottom-heavy then this contribution will take even longer with solar mass stars taking billions of years to start their dust production. Therefore, the majority of the stars could not aid in dust production at high redshifts. Past studies show that AGB stars in early galaxies need at least 300 Myr to dominate the dust production (DiCriscienzo2013; Valiante2017), which is too long to account for the earliest dusty galaxies.
@@
In recent decades, core-collapse supernovae (CCSNe) have been found to be a promising candidate for dust production in the early universe (e.g. Morgan2003; Dwek2007; Gall2011b; Gall2011c; Gall2011; Lesniewska2019; Pozzi2021; Dayal2022; Palla2024) because they originate from stars more massive than 8 M, which have lifetimes of only a few to several tens of million years. Observations in the local universe have shown that, within the first few decades after the SN explosion, 1 M of dust can be formed in the SNR ejecta (e.g. Lucy1989; Dunne2009; Matsuura2011; Gomez2012; Temim2012; Bevan2016; DeLooze2017; DeLooze2019; Niculescu-Duvaz2022). However, SNRs do not only produce dust but are also able to destroy a large amount of pre-existent dust due to their high-velocity forward shock. This shock induces gas-grain (sputtering) and grain-grain collisions (shattering), reducing the ISM dust grain sizes or destroying them completely (e.g. Barlow1978; Draine1979; Jones1994; Jones1996). Furthermore, the freshly produced ejecta dust needs to survive the reverse shock that forms in a subsample of SNRs with heavy pre-SN mass-loss or high density environment as a result of pressure build-up behind the forward shock. This could result in a significantly lower dust contribution from SNRs to the ISM than the commonly assumed 1 M (e.g. Kirchschlager2024b). When forward and reverse shock destruction are taken into account, it becomes unclear whether SNe can indeed make a net contribution to the dust budget.
@@
To resolve this uncertainty, it is crucial to constrain the dust destruction rate of the SNR forward shock, which encounters several tens of solar masses of dust during its evolution (e.g. Kirchschlager2022; Kirchschlager2024a). Many analytical and numerical forward shock dust destruction studies have been carried out (e.g. Jones1994; Jones1996; Nozawa2006; Bocchio2014; Slavin2015; Hu2019; Martinez-Gonzalez2019; Kirchschlager2022; Kirchschlager2024a; Vasiliev2024; Dedikov2025) and predict a wide range between almost total survival (Martinez-Gonzalez2019, 0.03% destruction rate) to total destruction (e.g. Bocchio2014, up to 100% destruction rate). Thus, SNRs can act as dust producers or destroyers depending on the considered forward shock study and the similarly uncertain reverse shock destruction rate (e.g. Bianchi2007; Nozawa2010; Silvia2010; Silvia2012; Bocchio2016; Micelotta2016; Slavin2020; Kirchschlager2019; Kirchschlager2023; Kirchschlager2024b; Vasiliev2024; Otaki2025).
@@
This large disparity does not only originate from different initial conditions but also from the various ways of implementing dust into codes (Slavin2024) and the considered dust transport and destruction mechanisms. Especially grain-grain collisions were expected to only have a small impact on the dust destruction and thus often neglected (e.g. Nozawa2006; Bianchi2007; Nozawa2010; Silvia2010; Silvia2012; Micelotta2016; Hu2019; Martinez-Gonzalez2019; Slavin2020; Vasiliev2024; Dedikov2025). However, Slavin2015 and Kirchschlager2019 found that the redistribution of grain sizes can work synergistically with other dust destruction mechanisms and thus lead to large destroyed dust masses. Furthermore, a magnetic field was also often not taken into account (e.g. Nozawa2006; Bianchi2007; Nozawa2010; Silvia2010; Silvia2012; Martinez-Gonzalez2019; Kirchschlager2020; Slavin2020; Kirchschlager2022; Vasiliev2024; Dedikov2025) although dust grains are charged and thus influenced by it. It was found that the effects of the Lorentz force can significantly change the dust destruction efficiency in SNRs provided that high enough magnetic field strengths are present (e.g. Shull1978; Draine1979; Fry2020; Kirchschlager2023; Kirchschlager2024a).
@@
In addition to the dust destruction mechanisms considered, the properties of the ambient medium surrounding the SN are expected to affect how much dust is destroyed by the forward shock. Most works consider a simple homogeneous ISM in which the SNR forward shock evolves spherically. There are just a few studies we know of that include an inhomogeneous ISM (Hu2019; Kirchschlager2022; Kirchschlager2024a; Vasiliev2024; Dedikov2025), in which the SNR evolution becomes non-spherical because the forward shock is slowed down significantly in the dense, cold filaments and can travel faster through voids. Since kinetic sputtering mainly depends on the shock velocity and thermal sputtering on the temperature, it is theorized that the dense and dust-rich filaments can shield dust from destruction. Simple small density fluctuations as considered by Kirchschlager2022 and Vasiliev2024 show no agreement as the former finds a slightly higher dust destruction compared to the homogeneous case and the latter does not provide such a comparison. A more comprehensive study of different density inhomogeneities of Dedikov2025 draws a clearer picture. They find that a clumpy ISM can reduce the dust destruction by up to $\sim$20%. More complex turbulence as studied by Hu2019 and Kirchschlager2024a shows the same trend. These studies drive their turbulence by setting off SNe every 1–2 Myr at random locations in an initially homogeneous medium and letting the system evolve for 0.4 Gyr or to a statistical steady state, respectively. Since the turbulence is only driven by SNe, it lacks small-scale structures that influence the SNR evolution, particularly in the early stages, and are therefore expected to affect the dust destruction by the forward shock. Additionally, other feedback mechanisms such as winds and photoionization, or gravity, which would affect the turbulent energy cascade, are neglected. Thus, although this type of turbulence ensures a purely physical origin, it often has different statistical properties compared to the observed multi-scale turbulence (e.g. Federrath2010) whose main driving mechanisms are still under debate (e.g. Miesch1994; MacLow2004; Brunt2009; Burkhart2019; Sartorio2021).
@@
Turbulence can also be driven in simulations by a forcing term which can be employed by an Ornstein-Uhlenbeck process to insert Fourier waves with random propagation directions, and at random locations and scales that go down to the Jeans length. These shocks then interact until the density and velocity probability density functions (PDFs) of the fluid quantities converge. Since small and large energy waves are injected, different kinds of modes will occur in the resulting turbulence. With this method, the initial conditions can be finetuned to closely resemble the observed turbulence with simulations.
@@
We use this approach to drive purely solenoidal turbulence in a highly resolved 3D simulation box, in order to resemble the density distribution seen in regions of the ISM, including small scale structures. Afterwards, an SN is injected into the driven inhomogeneities and evolved for 10 kyr. During this period, the SNR forward shock is the most destructive due to its high initial kinetic and thermal energy. Furthermore, it already encounters a significant amount of dust due to its rapid expansion. After 10 kyr, the SNR will continue expanding and destroying dust, however, with a lower destruction rate. By including thermal and kinetic sputtering, grain-grain collisions, and a magnetic field, we ensure that all known relevant dust transport and destruction mechanisms are considered. Thereby, we aim to improve constraints on the forward shock dust destruction efficiency during the most destructive phase of SNR evolution.
@@
We describe the methods in greater detail in Sect. 2, followed by Sect. 3 which shows and analyzes the results of our magnetohydrodynamical (MHD) and dust simulations. The findings are discussed and compared to previous studies in Sect. 4 and the paper is finally concluded in Sect. 5.
@@ 2 Methods
In this study, we perform 3D hydrodynamic simulations with the widely used quasi-Lagrangian code Arepo (Springel2010) which uses a moving mesh based on Voronoi tessellation. We drive turbulence using an Ornstein-Uhlenbeck process (Sect. 2.1), and consider different average ISM densities of $1\,\mathrm{cm^{-3}}$ (low density), $10\,\mathrm{cm^{-3}}$ (intermediate density), or $100\,\mathrm{cm^{-3}}$ (high density) and a turbulent Mach number of 0 (homogeneous), 0.3 (subsonic), 1 (transonic), or 3 (supersonic). Afterwards, an SN explosion is injected as a large amount of thermal energy, distributed over a few cells at the center of our simulation domain. The SNR evolution is followed for 10 kyr in 3D MHD simulations with the same code (Sect. 2.2). Since Arepo does not include detailed dust transport and destruction mechanisms, we use the dust post-processing code Paperboats (Kirchschlager2019) that assumes an initial dust distribution in the turbulent box and then calculates the dust transport and destruction based on gas and plasma drag, Lorentz acceleration, thermal and kinetic sputtering, grain-grain collisions, and accretion (Sect. 2.3). Using this workflow, we can estimate the dust destruction efficiency of SNR forward shocks in different ISM environments.
@@ 2.1 Turbulence driving
The 3D turbulence simulations are carried out purely hydrodynamically by using the turbulence driving within Arepo (Mocz2017; Mocz2018), which makes use of the Ornstein-Uhlenbeck process. The simulations start with a homogeneous medium in which the turbulence is driven until it is in a statistically stationary state, thus, until the density and velocity PDF, and the power spectrum converge. This method of driving lets us generate turbulence that closely resembles the observed ISM, as a visual comparison in Fig. 1 shows. There, the dust column density of one of our driven turbulence snapshots is qualitatively compared to a dust map of the Polaris flare from Herschel/SPIRE observations (Miville-Deschenes2010) [^1] We downloaded and visualized the level 2.5 data products from the Herschel Science Archive. to show that the morphological features and scales are similar.
@@
In the turbulence simulations, the mass of every Voronoi cell is kept approximately the same, which induces refinement but does not change the total amount of cells. We did not include additional refinement criteria for the turbulence driving. The box itself is 20 or 40 pc large in each coordinate direction, depending on the used ISM density and turbulence strength. These box sizes have been found to be large enough so that the SNR can evolve for up to 10 kyr without reaching the boundaries in every simulation. The boundaries are chosen to be periodic to ensure that no dust escapes from the box in post-processing. To be able to resolve filaments in the ISM that were observed to have a width of $\sim$0.1 pc (Arzoumanian2011), we use $512^{3}$ cells in total. Thus, the resolution in filaments is up to 0.008 pc (0.016 pc) whereas in voids it is 0.15 pc (0.3 pc) for the 20 pc (40 pc) box.
@@
The studied parameter sets include different turbulence strengths, defined by the turbulent Mach number $\mathcal{M}=\sigma_{v}/c$ that is dependent on the velocity dispersion $\sigma_{v}$ and the sound speed $c$ (Beattie2019). This Mach number determines the statistical distribution of the density and velocity, as can be seen in Fig. 2 with the density PDF. Lower turbulent Mach numbers only show slight density variations (e.g. $\mathcal{M}=0.3$, blue line), whereas higher turbulent Mach numbers result in dense filaments and empty voids (e.g. $\mathcal{M}=3$, pink line). Because the turbulent Mach number of the interstellar and circumstellar medium around stars can differ significantly depending on the exact location of the star, we take different Mach numbers into account – from subsonic ($\mathcal{M}=0.3$) to transonic ($\mathcal{M}=1$) and supersonic ($\mathcal{M}=3$), and compare them to the homogeneous case ($\mathcal{M}=0$). This Mach number range reflects typical conditions in the ISM, with observed values ranging from subsonic (a few tenths) to supersonic (e.g. Heyer2006; Ginsburg2013; Schneider2013; Marchal2021, $\mathcal{M}=0.17$–8).
@@
Figure 2: Density PDFs of the different turbulent Mach numbers $\mathcal{M}$ in the converged low density (1 cm-3) turbulence simulations. With a higher turbulent Mach number, cell densities ($\rho$) can have a larger deviation from the average density $\rho$¿.
@@
We also include different ISM densities since SNe can occur in media of various densities. Stars can remain in their birth molecular cloud throughout their lifetime, where they can explode close to dense filaments or in a medium that was emptied by previous SNRs, or drift away from the cloud to lower density media (MaykerChen2023; Sarbadhicary2023). Thus, we chose to study different average ISM number densities of $n=1,10\text{ and }100\,\mathrm{cm^{-3}}$ in addition to the set of studied turbulent Mach numbers to comprehensively cover various realistic SNR surroundings. The column density of such a turbulence simulation box can be seen in Fig. 1 (left). Furthermore, slices of all simulations that were carried out in this study are visualized in Appendix A, in Figs. A, 9, and A for the low, intermediate, and high density environments, respectively.
@@
The initial ISM temperature in these simulations is chosen to be $10^{4}\,\mathrm{K}$. In the interest of maintaining a smaller computational requirement, we do not include magnetic fields, cooling or self-gravity in the turbulence driving. We do include the first two a posteriori in the SN explosion simulations, as laid out later in this section.

### v2

@@ Abstract
Context. While supernova remnants (SNRs) have been observed to produce up to 1 M of dust, the amount of dust destroyed by the forward shock is poorly constrained, raising the question of whether they are net dust producers or destroyers.
Aims. Our aim was to estimate the dust destruction efficiency of SNR forward shocks in a realistically turbulent interstellar medium (ISM) during their most destructive phase, and to assess dust shielding by high-density filaments during this period.
Methods. We ran 3D high-resolution turbulence simulations for different turbulent Mach numbers (0–3) and average ISM densities (1–100 cm-3) to resemble observations of the turbulent ISM. We then set off a supernova explosion to trace its 3D magnetohydrodynamic evolution for the first 10 kyr. Finally, we ran post-processing simulations to investigate the dust transport and destruction by the SNR forward shock, taking into account gas and plasma drag, kinetic and thermal sputtering, and grain-grain collisions, and considering either silicates or carbonaceous dust.
Results. The dust destruction rate of the forward shock strongly depends on the average ISM density and turbulence strength, varying in the range 27–92% (0.85–11.0 M) in the studied 10 kyr. Overall, dust is less efficiently destroyed in a low-density medium ($1\,\mathrm{cm^{-3}}$, 27–57%) than in intermediate-density ($10\,\mathrm{cm^{-3}}$, 46–92%) and high-density ($100\,\mathrm{cm^{-3}}$, 73–87%) cases. The forward shock is found to destroy 8–34% less dust in high Mach turbulence compared to a homogeneous medium. Furthermore, carbonaceous grains are up to 21% more robust than silicates.
Conclusions. Filaments can partly shield dust from destruction in the first 10 kyr; however, always more than 0.85 M of dust is destroyed, making most SNRs dust sinks under the conditions explored in this work. The destruction efficiency of the SNRs with less than 1 M of destroyed dust has not yet plateaued, so that they are most likely also net dust destroyers by the end of their lifetimes.
@@ 1 Introduction
Understanding dust formation and destruction is key for planetary, stellar, and (extra-)galactic astronomy alike (cf. Draine 2003; Micelotta et al. 2018). Asymptotic giant branch (AGB) stars and supernova remnants (SNRs) are known to provide favorable conditions for dust production. Formed dust grains are also hypothesized to be able to grow efficiently in the interstellar medium (ISM; e.g., Draine 2009; Hirashita 2012; Mattsson and Andersen 2012; Zhukovska et al. 2016, 2018), significantly increasing the ISM dust budget over time. Even so, the relative importance of these dust formation and growth mechanisms across redshifts and the detailed underlying physical processes remain uncertain (cf. Schneider and Maiolino 2024).
@@
Asymptotic giant branch stars play an important role in the dust budget of local galaxies (e.g., Matsuura et al. 2009; Olofsson et al. 2010; Schneider et al. 2014; Dell’Agli et al. 2015). However, at high redshifts, where several galaxies have been observed to contain high dust masses (e.g., Pettini et al. 1994; Bertoldi et al. 2003; Watson et al. 2015; Bowler et al. 2022; Sommovigo et al. 2022; Akins et al. 2023; Witstok et al. 2023), the dust needs to be produced at a much faster rate than typical AGB dust production channels can provide. AGB stars are the end evolutionary stages of low- and intermediate-mass stars with masses of less than 8 M, and thus take several tens of million years to start contributing to the dust budget (e.g., Vassiliadis and Wood 1993; Valiante et al. 2009). If the initial mass function is bottom-heavy then this contribution will take even longer with solar mass stars taking billions of years to start their dust production. Therefore, the majority of the stars could not aid in dust production at high redshifts. Past studies show that AGB stars in early galaxies need at least 300 Myr to dominate the dust production (Di Criscienzo et al. 2013; Valiante et al. 2017), which is too long to account for the earliest dusty galaxies.
@@
In recent decades, core-collapse supernovae (CCSNe) have been found to be a promising candidate for dust production in the early universe (e.g., Morgan and Edmunds 2003; Dwek et al. 2007; Gall et al. 2011a, b, c; Leśniewska and Michałowski 2019; Pozzi et al. 2021; Dayal et al. 2022; Palla et al. 2024) because they originate from stars more massive than 8 M, which have lifetimes of only a few to several tens of million years. Observations in the local universe have shown that, within the first few decades after the SN explosion, 1 M of dust can be formed in the SNR ejecta (e.g., Lucy et al. 1989; Dunne et al. 2009; Matsuura et al. 2011; Gomez et al. 2012; Temim et al. 2012; Bevan and Barlow 2016; De Looze et al. 2017, 2019; Niculescu-Duvaz et al. 2022). However, SNRs do not only produce dust; they are also able to destroy a large amount of pre-existent dust, due to their high-velocity forward shock (e.g., Lakićević et al. 2015; Priestley et al. 2021, 2022). This shock induces gas-grain collisions (sputtering) and grain-grain collisions (shattering), thus reducing the ISM dust grain sizes or destroying them completely (e.g., Barlow 1978; Draine and Salpeter 1979; Jones et al. 1994, 1996). Furthermore, the freshly produced ejecta dust needs to survive the reverse shock that forms in a subsample of SNRs with heavy pre-SN mass-loss or high-density environment as a result of pressure build-up behind the forward shock. This could result in a significantly lower dust contribution from SNRs to the ISM than the commonly assumed 1 M (e.g., Kirchschlager et al. 2024b). When forward and reverse shock destruction are taken into account, it becomes unclear whether SNe can indeed make a net contribution to the dust budget.
@@
To resolve this uncertainty, it is crucial to constrain the dust destruction rate of the SNR forward shock, which encounters several tens of solar masses of dust during its evolution (e.g., Kirchschlager et al. 2022, 2024a). Many analytical and numerical forward shock dust destruction studies have been carried out (e.g., Jones et al. 1994, 1996; Nozawa et al. 2006; Bocchio et al. 2014; Slavin et al. 2015; Hu et al. 2019; Martínez-González et al. 2019; Kirchschlager et al. 2022, 2024a; Vasiliev and Shchekinov 2024; Dedikov and Vasiliev 2025) and predict a wide range of outcomes between almost total survival (Martínez-González et al. 2019, 0.03% destruction rate) to total destruction (e.g., Bocchio et al. 2014, up to 100% destruction rate). Thus, SNRs can act as dust producers or destroyers depending on the considered forward shock study and the similarly uncertain reverse shock destruction rate (e.g., Bianchi and Schneider 2007; Nozawa et al. 2010; Silvia et al. 2010, 2012; Bocchio et al. 2016; Micelotta et al. 2016; Slavin et al. 2020; Kirchschlager et al. 2019, 2023, 2024b; Vasiliev and Shchekinov 2024; Otaki et al. 2025).
@@
This large disparity originates from the different initial conditions, and also from the various ways of implementing dust into codes (Slavin 2024) and the considered dust transport and destruction mechanisms. In particular, grain-grain collisions were expected to only have a small impact on the dust destruction, and thus were often neglected (e.g., Nozawa et al. 2006; Bianchi and Schneider 2007; Nozawa et al. 2010; Silvia et al. 2010, 2012; Micelotta et al. 2016; Hu et al. 2019; Martínez-González et al. 2019; Slavin et al. 2020; Vasiliev and Shchekinov 2024; Dedikov and Vasiliev 2025). However, Slavin et al. (2015) and Kirchschlager et al. (2019) found that the redistribution of grain sizes can work synergistically with other dust destruction mechanisms, and thus lead to large amounts of destroyed dust masses. Furthermore, a magnetic field was also often not taken into account (e.g., Nozawa et al. 2006; Bianchi and Schneider 2007; Nozawa et al. 2010; Silvia et al. 2010, 2012; Martínez-González et al. 2019; Kirchschlager et al. 2020; Slavin et al. 2020; Kirchschlager et al. 2022; Vasiliev and Shchekinov 2024; Dedikov and Vasiliev 2025), although dust grains are charged and thus influenced by it. It was found that the effects of the Lorentz force can significantly change the dust destruction efficiency in SNRs provided that high enough magnetic field strengths are present (e.g., Shull 1978; Draine and Salpeter 1979; Fry et al. 2020; Kirchschlager et al. 2023, 2024a).
@@
In addition to the dust destruction mechanisms considered, the properties of the ambient medium surrounding the SN are expected to affect how much dust is destroyed by the forward shock. Most works consider a simple homogeneous ISM in which the SNR forward shock evolves spherically. There are just a few studies we know of that include an inhomogeneous ISM (Hu et al. 2019; Kirchschlager et al. 2022, 2024a; Vasiliev and Shchekinov 2024; Dedikov and Vasiliev 2025), in which the SNR evolution becomes non-spherical because the forward shock is slowed down significantly in the dense cold filaments and can travel faster through voids. Since kinetic sputtering mainly depends on the shock velocity and thermal sputtering on the temperature, it is theorized that the dense and dust-rich filaments can shield dust from destruction. Simple small density fluctuations, as considered by Kirchschlager et al. (2022) and Vasiliev and Shchekinov (2024), show no agreement as the former finds a slightly higher dust destruction compared to the homogeneous case and the latter does not provide such a comparison. A more comprehensive study of different density inhomogeneities of Dedikov and Vasiliev (2025) draws a clearer picture. They find that a clumpy ISM can reduce the dust destruction by up to $\sim$20%. More complex turbulence, as studied by Hu et al. (2019) and Kirchschlager et al. (2024a), shows the same trend. These studies drive their turbulence by setting off SNe every 1–2 Myr at random locations in an initially homogeneous medium and letting the system evolve for 0.4 Gyr or to a statistical steady state, respectively. Since the turbulence is only driven by SNe, it lacks small-scale structures that influence the SNR evolution, particularly in the early stages, and are therefore expected to affect the dust destruction by the forward shock. Additionally, other feedback mechanisms such as winds and photoionization, or gravity, which would affect the turbulent energy cascade, are neglected. Thus, although this type of turbulence ensures a purely physical origin, it often has different statistical properties compared to the observed multi-scale turbulence (e.g., Federrath et al. 2010) whose main driving mechanisms are still under debate (e.g., Miesch and Bally 1994; Mac Low and Klessen 2004; Brunt et al. 2009; Burkhart and Mocz 2019; Sartorio et al. 2021).
@@
Turbulence can also be driven in simulations by a forcing term which can be employed by an Ornstein-Uhlenbeck process to insert Fourier waves with random propagation directions, and at random locations and scales that go down to the Jeans length. These shocks then interact until the density and velocity probability density functions (PDFs) of the fluid quantities converge. Since small and large energy waves are injected, different kinds of modes will occur in the resulting turbulence. With this method, the initial conditions can be fine-tuned to closely resemble the observed turbulence with simulations.
@@
We used this approach to drive purely solenoidal turbulence in a highly resolved 3D simulation box, in order to resemble the density distribution seen in regions of the ISM, including small-scale structures. Afterward, an SN was injected into the driven inhomogeneities and evolved for 10 kyr. During this period, the SNR forward shock is the most destructive due to its high initial kinetic and thermal energy. Furthermore, it already encounters a significant amount of dust due to its rapid expansion. After 10 kyr, the SNR will continue expanding and destroying dust, but with a lower destruction rate. By including thermal and kinetic sputtering, grain-grain collisions, and a magnetic field, we ensured that all known relevant dust transport and destruction mechanisms were considered. In this way, we aim to improve constraints on the forward shock dust destruction efficiency during the most destructive phase of SNR evolution.
@@
We describe the methods in greater detail in Sect. 2, followed by Sect. 3, which shows and analyzes the results of our magnetohydrodynamic (MHD) and dust simulations. The findings are discussed and compared to previous studies in Sect. 4, and the paper concludes in Sect. 5.
@@ 2 Methods
For this study we performed 3D hydrodynamic simulations with the widely used quasi-Lagrangian code Arepo (Springel 2010), which uses a moving mesh based on Voronoi tessellation. We drove turbulence using an Ornstein-Uhlenbeck process (Sect. 2.1), and considered different average ISM densities of $1\,\mathrm{cm^{-3}}$ (low density), $10\,\mathrm{cm^{-3}}$ (intermediate density), or $100\,\mathrm{cm^{-3}}$ (high density) and a turbulent Mach number of 0 (homogeneous), 0.3 (subsonic), 1 (transonic), or 3 (supersonic). Afterward, an SN explosion was injected as a large amount of thermal energy, distributed over a few cells at the center of our simulation domain. The SNR evolution was followed for 10 kyr in 3D MHD simulations with the same code (Sect. 2.2). Since Arepo does not include detailed dust transport and destruction mechanisms, we used the dust post-processing code Paperboats (Kirchschlager et al. 2019) that assumes an initial dust distribution in the turbulent box and then calculates the dust transport and destruction based on gas and plasma drag, Lorentz acceleration, thermal and kinetic sputtering, grain-grain collisions, and accretion (Sect. 2.3). Using this workflow, we are able to estimate the dust destruction efficiency of SNR forward shocks in different ISM environments.
@@ 2.1 Turbulence driving
The 3D turbulence simulations were carried out purely hydrodynamically by using the turbulence driving within Arepo (Mocz et al. 2017; Mocz and Burkhart 2018), which makes use of the Ornstein-Uhlenbeck process. We started the simulations with a homogeneous medium in which we drove the turbulence until it was in a statistically stationary state, and thus until the density and velocity PDF, and the power spectra converged. This method of driving let us generate a turbulence that closely resembles the observed ISM, as a visual comparison in Fig. 1 shows. There, the dust column density of one of our driven turbulence snapshots is qualitatively compared to a dust map of the Polaris flare from Herschel/SPIRE observations (Miville-Deschênes et al. 2010) [^1] We downloaded and visualized the level 2.5 data products from the Herschel Science Archive. to show that the morphological features and scales are similar.
@@
In the turbulence simulations the mass of every Voronoi cell was kept approximately the same, which induced refinement but did not change the total number of cells. We did not include additional refinement criteria for the turbulence driving. The box itself was 20 or 40 pc large in each coordinate direction, depending on the used ISM density and turbulence strength. These box sizes were large enough so that the SNR could evolve for up to 10 kyr without reaching the boundaries in every simulation. The boundaries were chosen to be periodic to ensure that no dust escapes from the box in post-processing. To be able to resolve filaments in the ISM that are observed to have a width of $\sim$0.1 pc (Arzoumanian et al. 2011), we used $512^{3}$ cells in total. Thus, the resolution in filaments was up to 0.008 pc (0.016 pc) whereas in voids it was 0.15 pc (0.3 pc) for the 20 pc (40 pc) box.
@@
The studied parameter sets include different turbulence strengths, defined by the turbulent Mach number $\mathcal{M}=\sigma_{v}/c$ that is dependent on the velocity dispersion $\sigma_{v}$ and the sound speed $c$ (Beattie et al. 2019). This Mach number determines the statistical distribution of the density and velocity, as can be seen in Fig. 2 with the density PDF. Lower turbulent Mach numbers only show slight density variations (e.g., $\mathcal{M}=0.3$, blue line), whereas higher turbulent Mach numbers result in dense filaments and empty voids (e.g., $\mathcal{M}=3$, pink line). Because the turbulent Mach number of the interstellar and circumstellar medium around stars can differ significantly depending on the exact location of the star, we took different Mach numbers into account – from subsonic ($\mathcal{M}=0.3$) to transonic ($\mathcal{M}=1$) and supersonic ($\mathcal{M}=3$), and compared them to the homogeneous case ($\mathcal{M}=0$). This Mach number range reflects typical conditions in the ISM, with observed values ranging from subsonic (a few tenths) to supersonic (e.g., Heyer et al. 2006; Ginsburg et al. 2013; Schneider et al. 2013; Marchal and Miville-Deschênes 2021, $\mathcal{M}=0.17$–8).
@@
Figure 2: Density PDFs of the different turbulent Mach numbers $\mathcal{M}$ in the converged low-density (1 cm-3) turbulence simulations. With a higher turbulent Mach number, cell densities ($\rho$) can have a larger deviation from the average density $\rho$¿.
@@
We also included different ISM densities since SNe can occur in media of various densities. Stars can remain in their birth molecular cloud throughout their lifetime, where they can explode close to dense filaments or in a medium that is emptied by previous SNRs, or can drift away from the cloud to lower density media (Mayker Chen et al. 2023; Sarbadhicary et al. 2023). Thus, we chose to study different average ISM number densities of $n=1,10\text{ and }100\,\mathrm{cm^{-3}}$ in addition to the set of studied turbulent Mach numbers to comprehensively cover various realistic SNR surroundings. The column density of such a turbulence simulation box can be seen in Fig. 1 (left). Furthermore, slices of all simulations that were carried out in this study are visualized in Appendix A, in Figs. A, 9, and A for the low-, intermediate-, and high-density environments, respectively.
@@
The initial ISM temperature in these simulations was chosen to be $10^{4}\,\mathrm{K}$. In the interest of maintaining a smaller computational requirement, we did not include magnetic fields, cooling or self-gravity in the turbulence driving. We did include the first two a posteriori in the SN explosion simulations, as laid out later in this section.